Institutt for fysikk, NTNU FY1006/TFY4215 Innføring i kvantefysikk Vår 2012

Lecture notes in chemical physics

April 2012

These notes represent a summary of the lectures in Chemical physics:

1. Quantum mechanical calculations on molecules • The Born-Oppenheimer approximation

• The Hartree and the Hartree-Fock approximations

• The Hartree method

• The Hartree-Fock method

• LCAO: Linear Combination of Atomic Orbitals

• Geometry optimization, energy minimization

• Vibrational frequencies

• Chemical reactions and equilibria

• Quantum mechanical modeling of chemical reactions

2. Nomenclature in organic chemistry (not relevant for exam 2012)

3. Stereochemistry

Appendix: Molecular models from Spartan

Last updated: March 22 2012.